You are given the following information about an economy:
Gross Investment $40
Govt. purchases of goods & service 30
GNP 200
X – M - 20
Personal Tax 60
Govt. transfer 25
Interest payments from the Govt.
to domestic Pvt. Sector 15
Factor income received from the rest of the would 7
Factor payment made to rest of would 9
Calculate:
a. Consumption b) GDP c) Net factor payment from abroad
b. Pvt. Saving e) Public Saving.
1
Expert's answer
2017-05-05T09:39:08-0400
Gross Investment $40
Govt. purchases of goods & service(government spending) 30
GNP 200
X – M(net export ) - 20
Personal Tax 60
Govt. transfer 25
Interest payments from the Govt.
to domestic Pvt. Sector 15
Factor income received from the rest of the world 7
Factor payment made to rest of world 9
GDP=C+I+G+NX, where C-consumption spending, I-gross private domestic investment,
G-government spending, NX-net export (NX =X-M, X-export, M-import)
GDP=GNP-NFP,
NFP (net factor payments) = factor payments from abroad – factor payments to abroad.
1) NFP = 7-9= -2;
2) GDP = 200 – (-2) =200+2=202;
3) C = GDP-I-G-NX; C = 202 - 40 - 30 - (-20) =202-40-30+20=152;
Private savings = Y - T – C+TR+N
Private savings is defined as the total income (Y) (might be referred to as GDP) minus the personal tax (T) and consumption spending (C) plus Govt. transfer(TR) and Interest payments from the Govt.to domestic Pvt. Sector(N):
4) Private savings=202-60-152+25+15= 30;
Public Saving = (T – TR – N) – G.
5) Public saving = (60- 25- 15) – 30= -10
This suggests that the government is running a budget deficit.